OpenStudy (mitu12):

factor the polynomial correctly c^2-64

OpenStudy (mitu12):

@Aureyliant

OpenStudy (anonymous):

What do you think you should do first?

OpenStudy (mitu12):

i really have no clue on this one

OpenStudy (anonymous):

Do you know the difference of squares?

OpenStudy (mitu12):

i heard of it just forgot what it meant

OpenStudy (anonymous):

Trying to factor as a Difference of Squares : Theory : A difference of two perfect squares, A2 - B2 can be factored into (A+B) • (A-B) Proof : (A+B) • (A-B) = A^2 - AB + BA - B^2 = A^2 - AB + AB - B^2 = A^2 - B^2

OpenStudy (anonymous):

Make sense?

OpenStudy (mitu12):

not really

OpenStudy (anonymous):

Where?

OpenStudy (anonymous):

c^2 is a perfect square 64 is a perfect square

OpenStudy (anonymous):

Therefore, we apply the difference of squares.

OpenStudy (mitu12):

but how do we factor them

OpenStudy (anonymous):

Check : 64 is the square of 8 Check : c2 is the square of c1 Factorization is : (c + 8) • (c - 8)

OpenStudy (anonymous):

See? It's simple as that when I apply the difference of squares rules.

OpenStudy (mitu12):

so you basically just take the square root of eat number

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

refer to one of my first responses for the rules.

OpenStudy (mitu12):

so your answer will be c-8

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

I literally spoonfed you the answer.

OpenStudy (anonymous):

A^2 - B^2 can be factored into (A+B) • (A-B)

OpenStudy (mitu12):

so... (c-8)(c+8)

OpenStudy (anonymous):

Yes.

OpenStudy (mitu12):

thank you